OpenStudy (anonymous):

What is the slope represented by this table of values for a linear function? x - 2 , 4 ,6,8 y - 7,10, 13 ,16

OpenStudy (tamtoan):

do you learn slope formula and equation of a line yet?

OpenStudy (anonymous):

Yes but didnt get it all something to do with rise over run or something like that

OpenStudy (tamtoan):

well, if you already learn those...you only need 2 points to find the equation. and use other points to test out your result.. equation of a line is in the form y = mx + b where m is a slope and b is a constant. and you find m by this formula : m = (y2-y1)/(x2-x1) with 2 points (x1,y1) and (x2,y2) do you get this ? it's kind of long :)

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

sort of but still lost

OpenStudy (tamtoan):

ok let's find slope m first, pick 2 points (x1,y1) and (x2,y2) from your table there, tell me what is the slope m...with that formula ?

OpenStudy (anonymous):

okay so pick two numbers from the table (x-2 y-7 ) and then do what ?

OpenStudy (anonymous):

y values increase by 3 and x values increase by 2. slope = rise/run = 3/2

OpenStudy (tamtoan):

no, let's take the first 2 points, (x1,y1) = (2,7) and (x2,y2) = (4,10) ..now can you calculate the slope?

OpenStudy (anonymous):

now I remember how to do it the answer is 3/2 right ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you guys and tamtoan for showing me the process :)

OpenStudy (tamtoan):

np mathhater :) hcmathkim is a simpler method but you will probably need all those formulas soon :)